# CineTrans: Learning to Generate Videos with Cinematic Transitions via Masked Diffusion Models

### v1

@@ 1 Introduction
Endowed with extensive pre-training and sophisticated architectures, diffusion models (Blattmann et al., 2023; Ho et al., 2020; Rombach et al., 2022; Song et al., b) have demonstrated promising capabilities in generating videos with high visual quality and strong consistency comparable to real videos (Wan et al., 2025; Kong et al., 2024; Yang et al., 2024; Li et al., 2024). Particularly in the domain of text-to-video generation (T2V) (Wang et al., 2024; Yang et al., 2024; Li et al., 2023; Cho et al., 2024), the breakthrough has attracted considerable attention, highlighting the potential of diffusion models in mastering video creation (Wang et al., 2024; Xu et al., 2024; Chen et al., 2023a; 2024a). However, generating multi-shot videos with cinematic editing style and movie-like transitions from a brief user input continues to be a significant challenge.
@@
Existing work on long video generation can be broadly categorized into two aspects. First, larger models trained on massive datasets advance the capability of video interpretation and generation, enhancing the visual fidelity and maximum video length. (Kong et al., 2024; Wan et al., 2025; Yang et al., 2024). Second, techniques such as conditioning improve consistency across concatenated samples, enabling longer outputs through clip stitching (Zheng et al., 2024; He et al., 2023; Xie et al., 2024; Zhao et al., 2024). While both approaches partially support multi-shot sequence generation, they still exhibit notable limitations. Large-scale models primarily focus on single-shot videos due to the scarcity of shot transitions in their training datasets (Wang et al., 2023b; Ju et al., 2025; Wang et al., 2023a). Cinematic transitions are not guaranteed, let alone at precisely controlled positions. Moreover, the high computational cost and extended training time undermine the efficiency of this approach. Meanwhile, generating individual shots separately and concatenating them requires substantial manual intervention, and ignores prior knowledge from cinematic multi-shot dataset, resulting in cuts that often misalign with real-world editing styles. Additionally, many recent works often target narrow contexts, such as facial consistency (Zheng et al., 2024) or specific animated series (Dalal et al., 2025), which constrains their applicability to general video generation.
@@
While the aforementioned methods achieve certain multi-shot effects, very few explicitly focus on cinematic transitions within diffusion models. In this work, we propose CineTrans, a framework that produces multi-shot videos with cinematic transitions as shown in Figure 1. To support this, we develop a refined data processing pipeline that processes raw footage into a 250K video-text dataset. Our split-stitch procedure groups semantically related clips and removes gradual transitions, and we employ transition-aware models to annotate hierarchical captions. The resulting Cine250K provides frame-level shot labels and temporally-dense captions, preserving film editing style and making it well-suited for multi-shot video generation, while proving effective in enabling more natural shot transitions and stronger inter-shot consistency.
@@
  • We analyze attention maps in diffusion models for multi-shot video generation and observe a strong connection between attention probabilities and shot transitions. Building on this insight, we introduce a mask mechanism that enables cinematic transitions within diffusion models, leading to the CineTrans framework, which is effective even in a training-free setting.
@@ 2 Related Work
Diffusion-based Video Generation Diffusion-based approaches, built on the iterative denoising framework of Latent Diffusion Model (Rombach et al., 2022), utilize scaled-up model (Kong et al., 2024; Wan et al., 2025; Yang et al., 2024; Blattmann et al., 2023; Chen et al., 2023a; 2024a; He et al., 2022; Lin et al., 2024; Polyak et al., 2024) and large video datasets (Miech et al., 2019; Bain et al., 2021; Zellers et al., 2021; Xue et al., 2022; Wang et al., 2023a; b; Chen et al., 2024b; Xiong et al., 2025) to generate high-quality, prompt-adherent videos with extended durations. Due to their strong semantic understanding, certain pretrained models (Kong et al., 2024; Wan et al., 2025) can preliminarily generate multi-shot videos when provided with transition-specified prompts, but at the expense of vast computational resources, long training times, and unreliable, imprecise transitions. In contrast, our work offers frame-level stable cinematic transitions, seamlessly applied to the diffusion-based framework.
@@
Multi-Shot Video Generation. Recent work has explored multi-shot video generation, which can be categorized into two main approaches. The first generates each shot separately and then concatenates them, focusing on consistency between the generated shots. Animate-a-Story (He et al., 2023) uses motion structure retrieval for plot-aligned clips guided by text prompts. (Chen et al., 2023b) proposes a mask-based diffusion model to generate a smooth transition between shots. Zhuang et al. (2024) attempts to utilize the language model to generate prompts for different shots of a video. DreamFactory (Xie et al., 2024) employs an LLM-based framework with multi-agent collaboration and keyframe iteration design to ensure consistency and seamless transitions. MovieDreamer (Zhao et al., 2024) adopts a hierarchical autoregressive architecture for global coherence, and VGoT (Zheng et al., 2024) uses keyframes and identity-preserving embeddings to enforce temporal and character consistency. While these consistency-driven approaches are effective to some extent, they do not leverage real multi-shot video datasets and tend to overlook complex relationships between shots, including variations in camera angles and scenes, which are crucial in human-edited videos. The second approach modifies model architectures to generate multi-shot content directly. Test-Time training (Dalal et al., 2025) adds a layer for long multi-shot generation but is tied to a specific animated series, limiting generalization. Mask2DiT (Qi et al., 2025) applies shot-wise semantic masking yet focuses on text condition injection and assumes fixed shot durations. ShotAdapter (Kara et al., 2025) and LCT (Guo et al., 2025) introduce transition tokens and specialized positional encodings, respectively, yet exhibit low consistency or require large-scale training. In contrast, our method generates cinematic multi-shot videos in a single pass with flexible frame-level control, achieving strong performance even in a training-free setting and demonstrating generalizability across diverse scenarios.
3 Problem Formulation
@@
Diffusion models (Ho et al., 2020; Song et al., a) are a class of generative models learning to reverse a diffusion process, which learn to corrupt data via a predefined noisy Gaussian process and then invert that corruption through a trained neural network. Video diffusion models simply apply this same forward-reverse framework across temporal frames, yielding coherent video sequences $\mathbf{F}=\{f_{1},f_{2},\ldots,f_{N}\}$, where each element $f_{t}$ represents a video frame. A key component in these models is integrating the attention mechanism (Vaswani et al., 2017), which allows latent variables to focus on each other’s relevant information and can be formalized as:
@@
  • Consistency within and across shots. Within a shot, maintaining visual consistency is crucial for continuity. In contrast, inter-shot consistency emphasizes high-level semantic similarity over low-level details, i.e., maintaining consistency across shots despite substantial compositional differences. This means that its evaluation is not confined to specific attributes, such as composition or facial features, but is instead guided by film-editing conventions, ensuring applicability to general scenarios.
4 Dataset
First, the transition points are identified by Pyscenedetect (Castellano, 2024), resulting in fragmented segments. Adjacent segments with high similarity, measured by ImageBind (Girdhar et al., 2023) features, are then stitched together according to predefined rules to assemble an initial collection of 16M clips containing shot transitions. We filter this collection by aesthetic score, duration, and shot count to form the preliminary video dataset. Subsequently, since shot transitions can be categorized into instantaneous hard cuts and gradual changes that blur segment boundaries, we apply TransNetV2 (Soucek & Lokoc, 2024) to detect and remove all gradual transition frames which do not clearly belong to any single shot. This yields unambiguous segments with precise shot labels, including exact start and end frame indices. In the final step, each video receives both a general caption produced by LLaVA-Video-7b-Qwen2 (Zhang et al., 2024) for temporally-dense descriptions of cinematic transitions, and separate shot captions from LLaVA-NeXT (Liu et al., 2024).
Following the pipeline, Cine250K offers high‑aesthetic videos, precise shot labels, and hierarchical captions, thereby supplying rich prior knowledge for cinematic multi‑shot video generation and facilitating the production of videos with authentic film editing style. In Table 1, we compare Cine250K with other datasets. As a multi-shot video dataset with detailed shot labels, the high-quality content of Cine250K can significantly facilitate research and exploration in multi-shot video generation. The details are presented in Appendix A and B.
Table 1: Comparison of Cine250K and other video-text datasets.
InternVid (Wang et al., 2023b) 234M 11.7s 17.6 - High 720P
LVD-2M (Xiong et al., 2025) 2M 20.2s 88.7 - Medium Diverse
OpenVid-1M (Nan et al., 2024) 1M N/A N/A - High Diverse
Panda70M (Chen et al., 2024b) 70.8M 8.5s 13.2 Medium 720P
LLaVA-Video-178K (Zhang et al., 2024) 178K $\sim$ 40.4s $\sim$ 300 High Diverse
Shot2Story20K (Han et al., 2023) 20K 16s 201.8 Medium 720P
Shot2Story134K 134K N/A N/A Medium 720P
Cine250K (Ours) 250K 10.75s 148.79 High 720P
5 Methodology
In this section, we introduce our insight and method for the proposed CineTrans. Section 5.1 presents our observations in the attention maps of diffusion models, highlighting the differences between intra-shot and inter-shot correlations. Based on this observation, we introduce a block-diagonal mask mechanism in detail in Section 5.2 for cinematic multi-shot video generation. In Section 5.3, we discuss the implementation of inference. An overview of the methodology is shown in Figure 4.
@@ 5.1 Frame correlation in attention module
Recently, owing to the increased model size, dataset scale, and computational resources, some diffusion models (Kong et al., 2024; Wan et al., 2025) have demonstrated preliminary capabilities in generating multi-shot videos. However, how diffusion models internally model transitions within a video remains unclear and warrants further exploration. We hypothesize that, for correlation between two adjacent frames, there is a significant difference between transition points and non-transition points. The transition point involves a substantial shift, while a non-transition point requires continuity to maintain visual coherence, making them inherently divergent.
@@
We explore and visualize frame-wise attention maps in the context of the multi-shot video generation cases. As shown in Figure 4(a), it demonstrates strong correlations for intra-shot frames and weak correlations for inter-shot frames. More specifically, the attention probability matrix exhibits a block-diagonal structure, with each block corresponding to a shot. Figure 4(a) also illustrates the quantitative differences in attention probabilities across various layers of the diffusion models, highlighting the variations between intra-shot and inter-shot correlations. To further quantify this observation, we compute the average ratio of mean intra-shot to inter-shot attention probabilities (26.68) and assess its correspondence with the ground-truth shot boundaries via Pearson correlation, yielding r=0.71 (p<<0.01). This suggests the potential of leveraging the attention maps to guide the generation of multi-shot videos.
(a) A case of generated multi-shot video within diffusion model.
(b) The architecture of the mask mechanism.
@@ 5.2 Mask mechanism
The mask mechanism functions through two aspects. First, it aligns with the phenomenon observed in Section 5.1, conforming to diffusion models’ inherent understanding of cinematic multi-shot sequences. Second, unmasked layers enable full-frame token attention, allowing each token to attend to others across different shots, thereby establishing high-level semantic consistency. This facilitates the effective utilization of multi-shot video datasets, thereby generating multi-shot videos that align with film editing. We present the impact of applying the mask at different layers on the results in Appendix D.2.
@@ 5.3 Implementations
Customization. Building on our mask mechanism, we guide a diffusion model originally designed for single-shot generation to execute user-specified transitions, thus producing genuine multi-shot videos. For instance, by incorporating LoRA (Hu et al., 2022) weights, we achieve zero-shot generation of multi-shot sequences with enhanced consistency and user-defined styles or character attributes, even though those weights are originally trained on single‑shot videos.
(a) A sunset ocean panorama gradually shifts to a warm close‑up of sunlit clouds. (shot count: 3)
(b) A lone ash‑blond traveler with angular cheekbones walks through a dusk‑lit plaza. (shot count: 3)
@@ 6 Experiment
[…]
6.1 Implementation details
We implement CineTrans-Unet based on LaVie (Wang et al., 2024), and the mask mechanism is applied to the last six layers. We finetune the model on Cine250K using a total batch size of 128 for 20,000 steps. The learning rate is set to $1\times 10^{-4}$. CineTrans‑DiT extends Wan2.1‑T2V‑1.3B (Wan et al., 2025) by integrating a mask mechanism applied to transformer layers 7-28 and is released in two variants. The training‑free variant augments the block‑diagonal mask with the Visible-First‑Frame Attention described in Section 5.3, enabling the generation of high‑quality multi‑shot videos. The second variant applies LoRA fine-tuning (rank=64) on top of the former, using a total batch size of 256 for 2,800 steps. All experiments are conducted on NVIDIA A100 GPUs. Moreover, we apply LoRA weights to CineTrans-DiT on top of the training-free version to enable model customization, and the results are shown in Figure 6.
Figure 6: Training-free results of model customization using LoRA weights.
For baseline comparison, we select three categories: large‑scale T2V diffusion model, multi‑shot model, and customization model. CogVideoX1.5‑5B (Yang et al., 2024), HunyuanVideo Kong et al. (2024), and Wanx2.1‑T2V‑turbo [^2] https://tongyi.aliyun.com/wanxiang/ leverage large‑scale pretraining and thus possess strong semantic understanding. Each of these models is prompted with a general instruction specifying the desired shot count. StoryDiffusion (Zhou et al., 2024) first produces a sequence of semantically consistent images following both the general prompt and individual shot‑specific prompts, which CogVideoXI2V (Yang et al., 2024) then expands into a video. As a customization model, Cinematron [^3] https://civitai.com/api/download/models/1494601?type=Model&format=SafeTensor offers dedicated transition capabilities and employs the same sampling procedure as Hunyuan.

### v2

@@ 1 Introduction
Endowed with extensive pre-training and sophisticated architectures, diffusion models (stablevideodiffusion; ho2020denoising; rombach2022high; song2020score) have demonstrated promising capabilities in generating videos with high visual quality and strong consistency comparable to real videos (wan2025wan; kong2024hunyuanvideo; yang2024cogvideox; li2024longvideosurvey). Particularly in the domain of text-to-video generation (T2V) (wang2024lavie; yang2024cogvideox; li2023videogen; cho2024t2v), the breakthrough has attracted considerable attention, highlighting the potential of diffusion models in mastering video creation (wang2024lavie; xu2024easyanimate; chen2023videocrafter1; chen2024videocrafter2). However, generating multi-shot videos with cinematic editing style and movie-like transitions from a brief user input continues to be a significant challenge.
@@
Existing work on long video generation can be broadly categorized into two aspects. First, larger models trained on massive datasets advance the capability of video interpretation and generation, enhancing the visual fidelity and maximum video length. (kong2024hunyuanvideo; wan2025wan; yang2024cogvideox). Second, techniques such as conditioning improve consistency across concatenated samples, enabling longer outputs through clip stitching (videogenofthought; he2023animateastory; xie2024dreamfactory; zhao2024moviedreamer). While both approaches partially support multi-shot sequence generation, they still exhibit notable limitations. Large-scale models primarily focus on single-shot videos due to the scarcity of shot transitions in their training datasets (internvid; ju2024miradata; wang2023videofactory). Cinematic transitions are not guaranteed, let alone at precisely controlled positions. Moreover, the high computational cost and extended training time undermine the efficiency of this approach. Meanwhile, generating individual shots separately and concatenating them requires substantial manual intervention, and ignores prior knowledge from cinematic multi-shot dataset, resulting in cuts that often misalign with real-world editing styles. Additionally, many recent works often target narrow contexts, such as facial consistency (videogenofthought) or specific animated series (ttt), which constrains their applicability to general video generation.
@@
While the aforementioned methods achieve certain multi-shot effects, very few explicitly focus on cinematic transitions within diffusion models. In this work, we propose CineTrans, a framework that produces multi-shot videos with cinematic transitions as shown in Figure 1. To support this, we develop a refined data processing pipeline that processes raw footage into a dataset of 250K video-text pairs. Our split-stitch procedure groups semantically related clips and removes gradual transitions, and we employ transition-aware models to annotate hierarchical captions. The resulting Cine250K provides frame-level shot labels and temporally-dense captions, preserving film editing style and making it well-suited for multi-shot video generation, while proving effective in enabling more natural shot transitions and stronger inter-shot consistency.
@@
  • We analyze attention maps in diffusion models for multi-shot video generation and observe a strong connection between attention probabilities and shot transitions. Building on this insight, we introduce a mask mechanism that enables cinematic transitions within diffusion models, leading to the CineTrans framework, which is effective in a training-free setting.
@@ 2 Related Work
Diffusion-based Video Generation. Diffusion-based approaches, built on the iterative denoising framework of Latent Diffusion Model (rombach2022high), utilize scaled-up model (kong2024hunyuanvideo; wan2025wan; yang2024cogvideox; stablevideodiffusion; chen2023videocrafter1; chen2024videocrafter2; he2022latent; lin2024opensora; polyak2024moviegen) and large video datasets (howto100m; webvid; YT; HDVILA; wang2023videofactory; internvid; panda; lvd) to generate high-quality, prompt-adherent videos with extended durations. Owing to strong semantic capacity, certain pretrained models (kong2024hunyuanvideo; wan2025wan) can preliminarily generate multi-shot videos when provided with transition-specified prompts, but require vast computational resources, long training, and yield imprecise transitions. In contrast, our work offers frame-level stable cinematic transitions, seamlessly applied to the diffusion-based framework.
@@
Multi-Shot Video Generation. Recent work has explored multi-shot video generation, which can be categorized into two main approaches. The first generates each shot separately and then concatenates them, focusing on consistency between the generated shots. Animate-a-Story (he2023animateastory) uses motion structure retrieval for plot-aligned clips guided by text prompts. SEINE (chen2023seine) proposes a mask-based diffusion model to generate a smooth transition between shots. Vlogger (Zhuang_2024_CVPR) attempts to utilize the language model to generate prompts for different shots of a video. DreamFactory (xie2024dreamfactory) employs an LLM-based framework with multi-agent collaboration and keyframe iteration design. MovieDreamer (zhao2024moviedreamer) adopts a hierarchical autoregressive architecture for global coherence, and VGoT (videogenofthought) uses keyframes and identity-preserving embeddings to enforce temporal and character consistency. While these consistency-driven approaches are effective to some extent, they do not leverage real multi-shot video datasets and tend to overlook complex relationships between shots, including variations in camera angles and scenes, which are crucial in human-edited videos. The second approach modifies model architectures to generate multi-shot content directly. Test-Time training (ttt) adds a layer for long multi-shot generation but is tied to a specific animated series, limiting generalization. Mask2DiT (qi2025mask2dit) applies shot-wise semantic masking yet focuses on text injection and assumes fixed shot durations. ShotAdapter (kara2025shotadapter) introduces transition tokens but exhibits low consistency. LCT (guo2025longcontexttuning) enhances multi-shot understanding with specialized positional encodings but demands large-scale training. In contrast, our method generates cinematic multi-shot videos in a single pass with flexible frame-level control, achieving strong performance even in a training-free setting and demonstrating generalizability across diverse scenarios.
Temporal-Controlled Video Generation. Several recent works in video generation focus on controlling the temporal dimension, enabling the synthesis of longer videos with more dynamic and fine-grained temporal control cai2025ditctrl; li2025vstar; ouyang2024flexifilm. VSTAR li2025vstar draws inspiration from the band-matrix-like structure observed in the attention maps of real videos, and introduces a Temporal Attention Regularization strategy to enhance temporal dynamics and motion continuity during generation. Our targeted task, multi-shot video generation, can be regarded as a variant of temporally controlled video generation. It shifts the temporal control from frame-level semantic evolution to shot-level semantic transitions, which better aligns with real-world video editing practices and the narrative flow of long-form videos. Fundamentally, our proposed mask mechanism, like other temporally controlled generation methods, manipulates the attention distribution among visual tokens within the diffusion process to achieve more coherent and controllable video synthesis.
Masked Attention Mechanism. Masked attention has been widely used to regulate contextual dependencies and enable controllable generation. It has been applied in representation learning baobeit; tong2022videomae, generative modeling chang2022maskgit; yu2023magvit; luo2023notice, and fine-grained spatial or compositional control in diffusion models endo2024masked; wang2024compositional. Our work follows this line but employs attention masking for shot-level temporal control in video diffusion, ensuring coherence across shots while preserving intra-shot consistency.
3 Dataset
First, the transition points are identified by Pyscenedetect (pyscenedetect), resulting in fragmented segments. Adjacent segments with high similarity, measured by ImageBind (girdhar2023imagebind) features, are then stitched together according to predefined rules to assemble an initial collection of 16M clips containing shot transitions. We filter this collection by aesthetic score, duration, and shot count to form the preliminary video dataset. Subsequently, since shot transitions can be categorized into instantaneous hard cuts and gradual changes that blur segment boundaries, we apply TransNetV2 (soucek2024transnetv2) to detect and remove all gradual transition frames which do not clearly belong to any single shot. This yields unambiguous segments with precise shot labels, including exact start and end frame indices. In the final step, each video receives both a general caption produced by LLaVA-Video-7b-Qwen2 (zhang2024llavavideo) for temporally-dense descriptions of cinematic transitions, and separate shot captions from LLaVA-NeXT (liu2024llavanext).
Following the pipeline, Cine250K offers high‑aesthetic videos, precise shot labels, and hierarchical captions, thereby supplying rich prior knowledge for cinematic multi‑shot video generation and facilitating the production of videos with authentic film editing style. The details and comparison with previous video datasets are presented in Appendix A and B.
4 Methodology
(a) A case of generated multi-shot video within diffusion model.
(b) The architecture of the mask mechanism.
In this section, we introduce our insight and method for the proposed CineTrans. We first provide the necessary preliminaries in Section 4.1 to facilitate the following discussions. Section 4.2 presents our observations in the attention maps of diffusion models, highlighting the differences between intra-shot and inter-shot correlations. Based on this observation, we introduce a block-diagonal mask mechanism detailedly in Section 4.3 for cinematic multi-shot video generation. In Section 4.4, we discuss the implementation of inference. An overview of the methodology is shown in Figure 4.
4.1 Preliminary
@@
Diffusion models (ho2020denoising; song2020denoising) are a class of generative models learning to reverse a diffusion process, which learn to corrupt data via a predefined noisy Gaussian process and then invert that corruption through a trained neural network. Video diffusion models simply apply this same forward-reverse framework across temporal frames, yielding coherent video sequences $\mathbf{F}=\{f_{1},f_{2},\ldots,f_{N}\}$, where each element $f_{t}$ represents a video frame. A key component in these models is integrating the attention mechanism (vaswani2017attention), which allows latent variables to focus on each other’s relevant information and can be formalized as:
@@
  • Consistency within and across shots. Within a shot, visual consistency is crucial for continuity. In contrast, inter-shot consistency emphasizes high-level semantic similarity over low-level details, i.e., maintaining consistency across shots despite substantial compositional differences. This means that its evaluation is not confined to specific attributes, such as composition or facial features, but is guided by film-editing conventions, ensuring applicability to general scenarios.
@@ 4.2 Frame correlation in attention module
Recently, owing to the increased model size, dataset scale, and computational resources, some diffusion models (kong2024hunyuanvideo; wan2025wan) have demonstrated preliminary capabilities in generating multi-shot videos. However, how diffusion models internally model transitions within a video remains unclear and warrants further exploration. We hypothesize that the correlation between adjacent frames at transition points is significantly different from that at non-transition points. The transition point involves a substantial shift, while a non-transition point requires continuity to maintain visual coherence, making them inherently divergent.
@@
We explore and visualize frame-wise attention maps in the context of the multi-shot video generation cases. As shown in Figure 4(a), it demonstrates strong correlations for intra-shot frames and weak correlations for inter-shot frames. More specifically, the attention probability matrix exhibits a block-diagonal structure, with each block corresponding to a shot. Figure 4(a) also illustrates the quantitative differences in attention probabilities across various layers, highlighting the variations between intra-shot and inter-shot correlations. To quantify this observation, we compute the average ratio of mean intra-shot to inter-shot attention probabilities (26.68) and assess its correspondence with the ground-truth shot boundaries via Pearson correlation, yielding r=0.71 (p<<0.01). This suggests the potential of leveraging the attention maps to guide the generation of multi-shot videos.
@@ 4.3 Mask mechanism
The mask mechanism functions through two aspects. First, it aligns with the phenomenon observed in Section 4.2, conforming to diffusion models’ inherent understanding of cinematic multi-shot sequences. Second, unmasked layers enable full-frame token attention, allowing each token to attend to others across different shots, thereby establishing high-level semantic consistency. This facilitates the effective utilization of multi-shot video datasets, thereby generating multi-shot videos that align with film editing. We present the impact of masking different layers in Appendix D.2.
@@ 4.4 Implementations
Customization. Building on our mask mechanism, we guide a diffusion model originally designed for single-shot generation to execute user-specified transitions, thus producing genuine multi-shot videos. For instance, by incorporating LoRA (hu2022lora) weights, we achieve zero-shot generation of multi-shot sequences with enhanced consistency and user-defined styles or character attributes, even though those weights are originally trained on single‑shot videos.
@@ 5 Experiment
Figure 6: Results of transferring our method to the customization model. It enables multi-shot video generation while maintaining consistent character identity.
5.1 Implementation details
We implement CineTrans-Unet based on LaVie (wang2024lavie), and the mask mechanism is applied to the last six layers. We finetune the model on Cine250K with a batch size of 128 for 20,000 steps and the learning rate is $1\times 10^{-4}$. CineTrans‑DiT extends Wan2.1‑T2V‑1.3B (wan2025wan) by integrating the mask mechanism applied to transformer layers 7-28 and is released in two variants. The training-free variant augments the block-diagonal mask with the Visible-First-Frame Attention (Section 4.4). The second variant further applies LoRA fine-tuning (rank=64) with a batch size of 256 for 2,800 steps. All experiments are conducted on NVIDIA A100 GPUs. We also apply LoRA weights to CineTrans-DiT for model customization, as shown in Figure 6.
For baseline comparison, we select three categories: large‑scale T2V diffusion model, multi‑shot model, and customization model. CogVideoX1.5‑5B (yang2024cogvideox), HunyuanVideo (kong2024hunyuanvideo), and Wanx2.1‑T2V‑turbo [^2] https://tongyi.aliyun.com/wanxiang/ leverage large‑scale pretraining and thus possess strong semantic understanding. Each of these models is prompted with a general instruction specifying the desired shot count. StoryDiffusion (zhou2024storydiffusion) first produces a sequence of semantically consistent images following both the general prompt and individual shot‑specific prompts, which CogVideoXI2V (yang2024cogvideox) then expands into a video. As a customization model, Cinematron [^3] https://civitai.com/api/download/models/1494601?type=Model&format=SafeTensor offers dedicated transition capabilities and employs the same sampling procedure as Hunyuan.
@@ 5.1 Implementation details
CineTrans-DiT 0.7003 0.7858 0.1552 0.7874 0.1901 0.9673 0.9775 0.6508 0.2109